OpenStudy (reyes_gabbs):

measure of angle 1 = x measure of angle 2 = x-6

OpenStudy (vanesaretana):

m∠3 = 180 -(x + x - 6 ) = 180 - 2x + 6 = 186 - 2x

jhonyy9 (jhonyy9):

vanesa how you know that this is a triangle problem,exercise or not ? not is specified

OpenStudy (mathmale):

You, @reyes_gabbs, need to share the instructions for this problem. What is your goal? Is there a picture to share?